ATLAS UK Cloud Support
                    
    
        
            
            →
            
        
        Europe/London
    
                
                
                    
                        
                            
    
    
        
            
                
                Zoom
            
            
                
    
        
        
            
        
    
                        
                    
                
                
                
                    Zoom
                        
                        
                            , 
                        
                    
                
            Description
            Meeting to be held via Zoom (https://ukri.zoom.us/j/97404730356)
Password protected (same as OPs Mtg)
Outstanding tickets
- 149752 UKI-NORTHGRID-LANCS-HEP less urgent assigned 2020-12-02 16:07:00 Failovers from University of Lancaster to CERN backup proxy
	
- Number of stale cvmfs observed (also at Glasgow)
 - geoip issues; might be related to Stratum 1 updates?
 - refresh cache may be best option
 
 - 149750 UKI-SOUTHGRID-RALPP less urgent in progress 2020-12-02 10:16:00 UKI-SOUTHGRID-RALPP: unable to connect to host
	
- Problems in FTS transfers for ATLAS (not other VOs). CLI TPC transfers appear ok.
 
 - 149738 UKI-NORTHGRID-LANCS-HEP less urgent in progress 2020-12-02 15:55:00 UKI-NORTHGRID-LANCS-HEP: deletion errors
	
- Poor raid card showing issues with many simultaneous interactions (deletions) causing crashing.
 - Down to last 25% of data from draning of the seriver.
 - Stop draining for today; but should expect some file losses.
 
 - 149705 UKI-SCOTGRID-ECDF less urgent in progress 2020-11-30 11:52:00 UKI-SCOTGRID-ECDF: Low transfer efficiency due to TRANSFER [70] TRANSFER an end-of-file was reached …
	
- Load on headnode from httpd processes
		
- From Matt; method to mitigate high mem usage at lancs for http implemented. Might be related issues.
 
 
 - Load on headnode from httpd processes
		
 - 149362 UKI-SOUTHGRID-RALPP urgent in progress 2020-11-19 10:11:00 ATLAS CE failures on UKI-SOUTHGRID-RALPP-heplnx207
	
- heplnx207 still in downtime (ended post-meeting)
 
 - 148342 UKI-SCOTGRID-GLASGOW less urgent in progress 2020-11-27 10:00:00 UKI-SCOTGRID-GLASGOW with transfer efficiency degraded and many failures
	
- Disk 40; being drained withing decom. Raid set says ok, FS not.
 - AC / cooling issues in DPM server room
 
 - 146651 RAL-LCG2 urgent on hold 2020-10-16 11:56:00 singularity and user NS setup at RAL
	
- on hold, working on underlying issues
 
 - 142329 UKI-SOUTHGRID-SUSX top priority on hold 2020-11-05 10:52:00 CentOS7 migration UKI-SOUTHGRID-SUSX
	
- Arc-ce issues; not reporting back to the monitoring sites
 - Communication issue ? GridFTP looks to be working
 - Can the BDII / LDAP be queried (from offsite?)
		
- Status information usually through the BDII.
 
 - To contact the arc-devs?
 - To try an LDAP search against BDII
 - Patrick to report back to TB support.
 
 
CPU
- 
	
RAL
 - 
	
Northgrid
 - 
	
London
 - 
	
SouthGrid
 - 
	
Scotgrid
 - 
	
Downtime for DPM; Problems with Chillers and AC. Effectively shut down for the moment.
- Some replacements needed.
 
 - 
	
Prod is in DC; which is fine
 
Other new issues
Ongoing issues
- 
	
CentOS7 - Sussex
 - 
	
TPC http
- RAL TPC-http FTS tests working by converting // to / in path.
 
 - 
	
Oxford Storageless tests
 - 
	
10GB link working
 - 
	
Arc config needed; Sam to send to Vip
 - 
	
ECDF unreliable storage
- Rob to update ticket
 
 - 
	
Glasgow LOCALGROUPDISK
- Sam to aim to create Ceph pool.
 
 
News round-table
- 
	
Vip
- Production squid server failover yesterday;
 - CPU efficiency looks a bit lower?
 - prmon to be added: https://github.com/HSF/prmon in monitoring for storageless tests.
 
 - 
	
Dan
- Possible downtime 1wk on the 14th.
		
- Storm moving ahead to centos7
 
 - Next year disruption expected in DC, dates to be determined.
 
 - Possible downtime 1wk on the 14th.
		
 - 
	
Matt
- NTR; prepare for lost files.
 
 - 
	
Peter
- Considering options for CRC shifter
		
- Soliciting for CRC shifts.
 
 
 - Considering options for CRC shifter
		
 - 
	
Sam
 - 
	
NTR
 - 
	
Gareth
- Continue to work on cooling issues
 
 - 
	
JW
- NTR
 
 - 
	
Patrick
- NTR
 
 
AOB
                    There are minutes attached to this event.
                    Show them.